OpenStudy (anonymous):

solve the inequality : x^3-2X^2-X+2 less than or equal to 0

OpenStudy (anonymous):

\[x^3-2x^2-x+2 \le 0\]\[x^2(x-2)-(x-2) \le 0\]\[(x-2)(x^2-1) \le 0\] x-2=0 and x^2-1=0 x=2 x^2=1 x= +/- 1 Sub these values into the original equation to determine which way the inequality goes. Then you should get \[x \le -1\]\[ 1\le x \le 2\]